# Gravitational wave inference of star cluster properties from intermediate-mass black hole mergers

### v1

@@ IV.2 Isothermal assumption
The assumption of isothermal conditions in $r_{\rm a}\lesssim r\lesssim r_{\rm h}$ is valid as long as $r_{\rm a}$ is smaller than the core radius $r_{\rm c}$. Otherwise, the core velocity dispersion within the cusp region is enhanced by the presence of the central BH. Based on numerical simulations and an energy balance between the flux through $r_{\rm h}$ and $r_{\rm c}$ in systems with a central BH, Reference Heggie et al. (2007) derives that $r_{\rm c}/r_{\rm h}\simeq 4.3(M_{\rm BH}/M_{\rm cl})^{3/4}$. Thus, $r_{\rm c}/r_{\rm a}\simeq 0.6(M_{\rm cl}/M_{\rm BH})^{1/4}$. Since we are interested in IMBHs with a mass no larger than $2000M_{\odot}$ and clusters with $r_{\rm h,0}>0.1\,\rm pc$, we estimate that the ratio $r_{\rm c}/r_{\rm a}$ does not fall short of unity. At the same time, it has a typical value of $\approx 1.9$ in the parameter space region of interest for this work. Therefore, the enhancement in $\sigma$ is not an issue, because we deal with relatively light IMBHs.
@@ V Conclusions
A Python implementation of our code for reproducibility of the results of this work is publicly available on GitHub at the URL https://github.com/Kkritos/Star_clusters_from_IMBH_mergers/tree/main.
@@ Acknowledgements.
We thank Itai Linial, Brian Metzger, Francesco Iacovelli, Digvijay Wadekar, Mark Cheung, and Yossef Zenati for useful discussions. K.K. is supported by the Onassis Foundation - Scholarship ID: F ZT 041-1/2023-2024. K.K., L.R., and E.B. are supported by NSF Grants No. AST-2307146, PHY-2207502, PHY-090003, and PHY-20043, by NASA Grants No. 20-LPS20-0011 and 21-ATP21-0010, by the John Templeton Foundation Grant 62840, by the Simons Foundation, and by the Italian Ministry of Foreign Affairs and International Cooperation Grant No. PGR01167. K.K.Y.N. is supported by a Miller Fellowship and by the Croucher Foundation. This work was carried out at the Advanced Research Computing at Hopkins (ARCH) core facility (rockfish.jhu.edu), which is supported by the NSF Grant No. OAC-1920103.
@@ Appendix A Tidal capture radius
Stars on parabolic orbits that approach a BH with a pericenter distance $r_{\rm p}>r_{\rm T}$ (where $r_{\rm T}$ is the tidal radius) are not tidally disrupted. However, they may still be tidally captured on more compact orbits around the BH as long as the total tidal energy dissipated during the first pericenter passage ($\Delta E_{\rm tid}$) exceeds the orbital energy at infinity ($T_{\infty}$). In this appendix, we estimate the critical radius $r_{\rm C}$ for tidal capture. In general, this parameter depends on the relative velocity between the BH and the star at infinity ($v_{\infty}$), as well as on the structure of the star (here modeled as a polytrope with polytropic index $n$). Since captured stars eventually reach $r_{\rm T}$ through tidal energy dissipation at each pericenter passage and the cross section for capture is larger than the TDE cross-section, the loss cone radius is effectively $r_{\rm C}$.
@@ Appendix B Cluster dissolution and BH feedback
In the isolated Hénon’s model, the ejection rate is low (about 0.74% of stars are ejected every relaxation time Ambartsumian (1985)), and the systems of interest do not evaporate within a Hubble time. Moreover, we have verified that the BH-to-cluster mass ratio in this scenario always remains below $4\%$ and typically at the sub-percent level, and the number of stars consumed by the BH is a tiny fraction of the total number. Thus the evolution of the global system approximately decouples from that of the BH, and the equations for $(dN_{\star}/dt,dr_{\rm h}/dt)$ and $d{M}_{\rm BH}/dt$ separate. This separation allows us to find a closed-form solution for the evolution of the system (notice that $dN_{\star}/dt$ and $dr_{\rm h}/dt$ are still coupled through the relaxation time). The growth of the BH strongly depends on the evolution of the cluster as a whole through the evolution of the velocity dispersion: see Eq. (1b).
@@
The feedback of the BH into the evolution of the global properties of the cluster enters through the contribution of the consumption rate in removing stars from the system, as well as the contribution of $M_{\rm BH}$ to $M_{\rm cl}$. In general, the full set of equations for the evolution of the system becomes:
@@ Appendix C Parameter estimation of IMBH binaries
We generate the GW signals using the IMRPhenomXPHM waveform model Pratten et al. (2021), a state-of-the-art quasi-circular model that includes both spin precession effects and higher-order modes. Higher-order modes are expected to be important for both detection Chandra et al. (2020) and PE Fairhurst et al. (2023) of systems with IMBH components. For each IMBH binary, we perform Bayesian PE on the whole set of 15 parameters that characterize the waveform model, namely,

### v2

@@ IV.2 Isothermal assumption
The assumption of isothermal conditions in $r_{\rm a}\lesssim r\lesssim r_{\rm h}$ is valid as long as $r_{\rm a}$ is smaller than the core radius $r_{\rm c}$. Otherwise, the core velocity dispersion within the cusp region is enhanced by the presence of the central BH. Based on numerical simulations and an energy balance between the flux through $r_{\rm h}$ and $r_{\rm c}$ in systems with a central BH, Ref. Heggie et al. (2007) derives that $r_{\rm c}/r_{\rm h}\simeq 4.3(M_{\rm BH}/M_{\rm cl})^{3/4}$. Thus, $r_{\rm c}/r_{\rm a}\simeq 0.6(M_{\rm cl}/M_{\rm BH})^{1/4}$. Since we are interested in IMBHs with a mass no larger than $2000M_{\odot}$ and clusters with $r_{\rm h,0}>0.1\,\rm pc$, we estimate that the ratio $r_{\rm c}/r_{\rm a}$ does not fall short of unity. At the same time, it has a typical value of $\approx 1.9$ in the parameter space region of interest for this work. Therefore, the enhancement in $\sigma$ is not an issue, because we deal with relatively light IMBHs.
@@ V Conclusions
A Python implementation of our code for reproducibility of the results of this work is publicly available on GitHub at the git .
@@ Acknowledgements.
We thank Itai Linial, Brian Metzger, Francesco Iacovelli, Digvijay Wadekar, Mark Cheung, and Yossef Zenati for useful discussions. K.K. is supported by the Onassis Foundation - Scholarship ID: F ZT 041-1/2023-2024. K.K., L.R., and E.B. are supported by NSF Grants No. AST-2307146, No. PHY-2207502, No. PHY-090003, and No. PHY-20043, by NASA Grants No. 20-LPS20-0011 and No. 21-ATP21-0010, by the John Templeton Foundation Grant 62840, by the Simons Foundation, and by the Italian Ministry of Foreign Affairs and International Cooperation Grant No. PGR01167. K.K.Y.N. is supported by a Miller Fellowship and by the Croucher Foundation. This work was carried out at the Advanced Research Computing at Hopkins (ARCH) core facility roc , which is supported by the NSF Grant No. OAC-1920103.
@@ Appendix A Tidal capture radius
Stars on parabolic orbits that approach a BH with a pericenter distance $r_{\rm p}>r_{\rm T}$ (where $r_{\rm T}$ is the tidal radius) are not tidally disrupted. However, they may still be tidally captured on more compact orbits around the BH as long as the total tidal energy dissipated during the first pericenter passage ($\Delta E_{\rm tid}$) exceeds the orbital energy at infinity ($T_{\infty}$). In this appendix, we estimate the critical radius $r_{\rm C}$ for tidal capture. In general, this parameter depends on the relative velocity between the BH and the star at infinity ($v_{\infty}$), as well as on the structure of the star (here modeled as a polytrope with polytropic index $n$). Since captured stars eventually reach $r_{\rm T}$ through tidal energy dissipation at each pericenter passage and the cross section for capture is larger than the TDE cross section, the loss cone radius is effectively $r_{\rm C}$.
@@ Appendix B Cluster dissolution and BH feedback
In the isolated Hénon’s model, the ejection rate is low (about 0.74% of stars are ejected every relaxation time Ambartsumian (1985)), and the systems of interest do not evaporate within a Hubble time. Moreover, we have verified that the BH-to-cluster mass ratio in this scenario always remains below $4\%$ and typically at the subpercent level, and the number of stars consumed by the BH is a tiny fraction of the total number. Thus the evolution of the global system approximately decouples from that of the BH, and the equations for $(dN_{\star}/dt,dr_{\rm h}/dt)$ and $d{M}_{\rm BH}/dt$ separate. This separation allows us to find a closed-form solution for the evolution of the system (notice that $dN_{\star}/dt$ and $dr_{\rm h}/dt$ are still coupled through the relaxation time). The growth of the BH strongly depends on the evolution of the cluster as a whole through the evolution of the velocity dispersion: see Eq. (1b).
@@
The feedback of the BH into the evolution of the global properties of the cluster enters through the contribution of the consumption rate in removing stars from the system, as well as the contribution of $M_{\rm BH}$ to $M_{\rm cl}$. In general, the full set of equations for the evolution of the system becomes
@@ Appendix C Parameter estimation of IMBH binaries
We generate the GW signals using the IMRPhenomXPHM waveform model Pratten et al. (2021), a state-of-the-art quasicircular model that includes both spin precession effects and higher-order modes. Higher-order modes are expected to be important for both detection Chandra et al. (2020) and PE Fairhurst et al. (2024) of systems with IMBH components. For each IMBH binary, we perform Bayesian PE on the whole set of 15 parameters that characterize the waveform model, namely,